A set of three scores consist of three values 3,7, and 2. (By Sigma I mean this symbol \sum).

Sigma2x-2=  

35, 33, 22, or 6?

Sigma(2x)2=

248, 124, 61, or 39?
Answer #1

1)here \sum2x-2 =(2*3+2*7+2*2)-2=22

2) \sum(2x)2   =(2*3)2+(2*7)2+(2*2)2 =248

( please revert for any clarification)